OpenStudy (anonymous):

Jeannie has some $10 bills and some $20 bills.if she has 273 bills worth a total of $4370,how many of the bills are $10 bills and how many are $20 bills.?

OpenStudy (***[isuru]***):

ok... If the amount of $20 bills r "x" then the amount of $10 bills r "(273 - X)" and if their sum is $4370... ya can say 20X + 10(273 - X) = 4370 now solve this for x did u get it ?

OpenStudy (anonymous):

Can you pls complete it

OpenStudy (***[isuru]***):

sure.. 20x + 2730 -10x = 4370 10x =1640 x = 164 soo... no. of $20 bills r 164 now.. no.of $10 bills r (273 - x) =109 hope this will help ya !!!

OpenStudy (anonymous):

not familiar with this equation

OpenStudy (anonymous):

is that final answer

OpenStudy (anonymous):

Solve for x LOL

OpenStudy (***[isuru]***):

yeah!!!

OpenStudy (***[isuru]***):

y ? do u still have a problem understanding that ? we could help ya if u be more specific to the point... I mean if u can say wt part is not clear ....

OpenStudy (anonymous):

I have never come across this type of word problem before thats why I need help in completing it then I can use it to practice on my own

OpenStudy (***[isuru]***):

Ok... so.. let's start it from the bottom.. do we know the no.of bills we have separately at the beginning ?

OpenStudy (***[isuru]***):

r u still there ?